OpenStudy (anonymous):

The six faces of a dice are numbered -3, -2, -1, 0, 1 and 2. If the dice is rolled twice and the two numbers are multiplied together, what is the probability that the result is a negative?

Parth (parthkohli):

When you multiply a negative with a positive, then the result is negative. So the probability of a negative number when they are multiplied is given by: P(Negative) * P(Positive)

Parth (parthkohli):

P(Negative) is 3/6 P(Non-negative) is 3/6 3/6 * 3/6 = 1/2 * 1/2 = 1/4

Parth (parthkohli):

So that is the final probability.

OpenStudy (anonymous):

P(negative no.) = 2* P(negative) *P(positive) :since your are having two dices P(negative) = 3/6 P(positive) = 2/6 therefore P(negative no.) = 1/3

Parth (parthkohli):

Oh yes